Part of an ANOVA table is shown below.

Source of
Variation

Sum of
Squares

Degrees
of Freedom

Mean
Square

F

Between Treatments

64

8

Within Treatments (Error)

2

Total

100

The number of degrees of freedom corresponding to between-treatments is

a.

3.

b.

4.

c.

2.

d.

18.
Answer #1

Given

Mean square within = MSSW = 2

F = 8

SS Between =SSB = 64

We know that

F = MSSB/MSSW

8 = MSSB/2

MSSB = mean square between = 16

We also know that

MSSB = SSB / DF OF SSB

We already find the value of MSSB

16 = 64/DF

DF OF BETWEEN TREATMENT = 64/16

DF OF BETWEEN TREATMENT = 4

Hence option (b) is correct.